OpenStudy (anonymous):

If g(x)=4x+3, then (gog)(x) is

jimthompson5910 (jim_thompson5910):

(gog)(x) is the same as g(g(x))

OpenStudy (anonymous):

how ? how do you solve

jimthompson5910 (jim_thompson5910):

g(x)=4x+3 g(g(x))=4( g(x) )+3 g(g(x))=4( 4x+3 )+3 keep going to simplify

OpenStudy (anonymous):

16x+12+3

jimthompson5910 (jim_thompson5910):

one more step

OpenStudy (anonymous):

16x+15

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

whats the next step

jimthompson5910 (jim_thompson5910):

that's it, you're done

jimthompson5910 (jim_thompson5910):

g(g(x)) = 16x + 15 so (gog)(x) = 16x + 15

OpenStudy (anonymous):

K thanks could you help with another problem similar

jimthompson5910 (jim_thompson5910):

sure